Use a proportion to solve each problem
The tax on a property with an assessed value of $65,000 is $720. Find the tax on a property with an assessed value of $162,500

Answer by checkley77(12844)

720/65,000=x/162,500
65,000x=720*162,500
65,000x=117,000,000
x=117,000,000/65,000
x=$1,800 ans. for the tax or the 162,500 property.